# Uncertainty Quantification with the Empirical Neural Tangent Kernel

### v1

@@ D.1 Comparison to SLU
Here we compare the performance of our method against the competing SLU method. See Appendix B for a discussion of the differences between NUQLS and SLU. Due to the extensive experimental details given in (Miani et al., 2024), we run certain experiments from this work and report the performance of NUQLS against the SLU results found in (Miani et al., 2024). In Table 3, we compare NUQLS against SLU on OoD detection using the AUC-ROC metric, on a smaller single-layer MLP and a larger LeNet model. The MLP is trained on the MNIST dataset, while the LeNet model is trained on the FashionMNIST dataset. For MNIST, the OoD datasets are FashionMNIST, KMNIST, and a Rotated MNIST dataset, where for each experiment run, we compute the average AUC-ROC score over a range of rotation angles $(15,30,45,60,75,90,105,120,135,150,165,180)$. For FashionMNIST, the OoD datasets are MNIST, and the average Rotated FashionMNIST dataset. The AUC-ROC metric was calculated using the variance of the logits, summed over the classes for each test point, as a score. We observe that NUQLS has a better AUC-ROC value over all OoD datasets, often by a significant margin.
@@
D.2 Comparison to Sampling-LLA, VaLLA and ELLA
We now compare NUQLS against Sampling-LLA, VaLLA and ELLA. Due to issues with convergence, performance, memory usage and package compatibility when either running source code or implementing methods from instructions given in (Antorán et al., 2022) and (Ortega et al., 2023), we instead compare NUQLS against the results for Sampling-LLA, VaLLA and ELLA taken verbatim from (Ortega et al., 2023, Figure 3). We train a $2$-layer MLP, with $200$ hidden-units in each layer, on the MNIST dataset, according to the experimental details given in (Ortega et al., 2023). We then compare the accuracy, NLL, ECE, Brier score and the OOD-AUC metric of NUQLS against those reported for Sampling-LLA, VaLLA and ELLA. The results are shown in Table 4. We display the mean and standard deviation for NUQLS; as is quoted in (Ortega et al., 2023), the standard deviation for the other methods was below $10^{-4}$ in magnitude, and was thus omitted. We see that NUQLS outperforms Sampling-LLA, VaLLA and ELLA in accuracy, NLL, ECE, and Brier score, and is within one-sixth of a standard deviation of the leading OOD-AUC value, attained by Sampling-LLA. While we cannot directly comment on differences in computation time between our method and these LLA extensions, we were able to run the source code for VaLLA for the experiments in Table 2, where NUQLS was an order-of-magnitude faster than VaLLA in wall-time. In (Ortega et al., 2023, Figure 3 (right)), we also observe that VaLLA is an order-of-magnitude faster than both Sampling-LLA and ELLA. We also compare with (Ortega et al., 2023, Table 1), where a ResNet20 model is trained on CIFAR-10. The results are displayed in Table 5. We see that NUQLS is competitive with all other methods in this setting.
@@ D.2 Comparison to Sampling-LLA, VaLLA and ELLA
Table 5: Comparing performance of NUQLS against Sampling-LLA, VaLLA and ELLA on CIFAR-10, with ResNet20. Original results taken from (Ortega et al., 2023).
Model ACC NLL ECE
ELLA $92.5$ $0.233$ $0.009$
Sampled LLA $92.5$ $0.231$ $\mathbf{0.006}$
VaLLA $\mathbf{92.6}$ $\mathbf{0.228}$ $0.007$
NUQLS $92.4\pm$$0.1$ $0.236\pm$$0.001$ $0.008\pm$$0.001$
D.3 Toy Regression BDE
Figure 4 displays the performance of BDE on the toy regression problem from Figure 1.
Figure 4: Comparison of BDE, DE and NUQLS on the toy regression problem from Figure 1. We can see that the uncertainty of the BDE method is quite small.
D.4 UCI Regression
We present the results for select UCI regression datasets in Table 1 in the main body; we present results for several more datasets in Table 6.
@@ D.4 UCI Regression
Concrete NUQLS $\mathbf{0.330\pm}$$\mathbf{0.047}$ $\mathbf{-0.316\pm}$$\mathbf{0.501}$ $\mathbf{0.003\pm}$$\mathbf{0.001}$ $\mathbf{0.185}$
DE $0.379\pm$$0.019$ $\mathbf{-0.574\pm}$$\mathbf{0.098}$ $\mathbf{0.002\pm}$$\mathbf{0.002}$ $29.047$
LLA $\mathbf{0.333\pm}$$\mathbf{0.050}$ $\mathbf{-0.294\pm}$$\mathbf{0.479}$ $\mathbf{0.003\pm}$$\mathbf{0.002}$ $0.297$
SWAG $\mathbf{0.334\pm}$$\mathbf{0.050}$ $\mathbf{-0.562\pm}$$\mathbf{0.224}$ $0.009\pm$$0.006$ $36.262$
Naval NUQLS $\mathbf{0.049\pm}$$\mathbf{0.012}$ $\mathbf{-2.546\pm}$$\mathbf{0.134}$ $\mathbf{0.002\pm}$$\mathbf{0.002}$ $\mathbf{0.295}$
DE $0.076\pm$$0.006$ $-1.761\pm$$0.250$ $0.093\pm$$0.040$ $96.570$
LLA $0.070\pm$$0.022$ $25.292\pm$$17.570$ $0.192\pm$$0.029$ $129.659$
SWAG $1.130\pm$$1.500$ $0.303\pm$$1.091$ $0.084\pm$$0.022$ $92.662$
CCPP NUQLS $0.244\pm$$0.008$ $-0.885\pm$$0.020$ $\mathbf{0.000\pm}$$\mathbf{0.000}$ $\mathbf{0.174}$
@@
LLA $0.243\pm$$0.007$ $29.420\pm$$4.565$ $0.163\pm$$0.008$ $32.048$
SWAG $0.252\pm$$0.012$ $-0.849\pm$$0.038$ $\mathbf{0.001\pm}$$\mathbf{0.002}$ $66.833$
Wine NUQLS $\mathbf{0.789\pm}$$\mathbf{0.042}$ $\mathbf{0.284\pm}$$\mathbf{0.066}$ $\mathbf{0.001\pm}$$\mathbf{0.000}$ $\mathbf{0.115}$
@@
LLA $\mathbf{0.792\pm}$$\mathbf{0.041}$ $1.012\pm$$0.182$ $0.009\pm$$0.004$ $0.340$
SWAG $\mathbf{0.798\pm}$$\mathbf{0.038}$ $\mathbf{0.367\pm}$$\mathbf{0.103}$ $0.005\pm$$0.003$ $11.807$
Yacht NUQLS $\mathbf{0.042\pm}$$\mathbf{0.013}$ $\mathbf{-1.561\pm}$$\mathbf{2.319}$ $\mathbf{0.012\pm}$$\mathbf{0.010}$ $\mathbf{0.164}$
DE $0.647\pm$$0.121$ $-2.032\pm$$0.349$ $\mathbf{0.016\pm}$$\mathbf{0.008}$ $40.132$
LLA $\mathbf{0.043\pm}$$\mathbf{0.014}$ $\mathbf{-2.733\pm}$$\mathbf{0.468}$ $\mathbf{0.011\pm}$$\mathbf{0.006}$ $0.177$
SWAG $\mathbf{0.044\pm}$$\mathbf{0.014}$ $-2.565\pm$$0.118$ $0.067\pm$$0.025$ $15.822$
Song NUQLS $\mathbf{0.839\pm}$$\mathbf{0.014}$ $0.646\pm$$0.056$ $0.001\pm$$0.000$ $\mathbf{91.673}$
DE $\mathbf{0.846\pm}$$\mathbf{0.006}$ $\mathbf{0.180\pm}$$\mathbf{0.013}$ $0.005\pm$$0.000$ $2562.789$
LLA-D $\mathbf{0.851\pm}$$\mathbf{0.029}$ $0.456\pm$$0.093$ $\mathbf{0.000\pm}$$\mathbf{0.000}$ $210.429$
SWAG $0.845\pm$$0.002$ $0.680\pm$$0.062$ $0.003\pm$$0.001$ $3274.440$
D.5 eNUQLS
In Figure 5 we demonstrate the performance of an ensembled version of NUQLS, eNUQLS. This method is similar to a Mixture of Laplace Approximations (Eschenhagen et al., 2021). We observe excellent separation of the variances between correct and incorrect/OoD test groups for eNUQLS, especially for CIFAR-10 on ResNet9. Note that there is significant cost to ensembling our method, and we provide this figure simply to illustrate performance capacity.
@@
D.6 Image Classification
The median and sample skew for the VMSP in Figure 2 is found in Table 7.
Table 7: Sample median and sample skew of variance from Figure 2. IDC = ID correct, IDIC = ID incorrect, FMNIST = FashionMNIST. The median and skew for a method is compared against a baseline method. We expect positive skew for the ID correct (IDC) variances, and a negative skew for the ID incorrect (IDIC) and OoD variances. If the median or skew for a method is worse than the corresponding baseline, then it is written in gray.
@@ D.6 Image Classification
ResNet9 IDC $\downarrow$ $2.45\times 10^{-8}$ $1.31\times 10^{-5}$ $4.64\times 10^{-7}$ $5.05\times 10^{-7}$ $9.02\times 10^{-10}$ $0.020$
@@
ResNet50 IDC $\downarrow$ $1.00\times 10^{-8}$ $8.73\times 10^{-4}$ $0.0449$ $1.12\times 10^{-4}$ $3.25\times 10^{-4}$ $0.020$
@@
Appendix E Experiment Details

### v2

@@ H.2 Comparison to SLU
Here we compare the performance of our method against the competing SLU method. See Appendix˜B for a discussion of the differences between NUQLS and SLU. Due to the extensive experimental details given in (Miani et al., 2024a), we run certain experiments from this work and report the performance of NUQLS against the SLU results found in (Miani et al., 2024a). In Table˜5, we compare NUQLS against SLU on OoD detection using the AUC-ROC metric, on a smaller single-layer MLP and a larger LeNet model. The MLP is trained on the MNIST dataset, while the LeNet model is trained on the FashionMNIST dataset. For MNIST, the OoD datasets are FashionMNIST, KMNIST, and a Rotated MNIST dataset, where for each experiment run, we compute the average AUC-ROC score over a range of rotation angles (15,30,45,60,75,90,105,120,135,150,165,180)(15,30,45,60,75,90,105,120,135,150,165,180). For FashionMNIST, the OoD datasets are MNIST, and the average Rotated FashionMNIST dataset. The AUC-ROC metric was calculated using the variance of the logits, summed over the classes for each test point, as a score. We observe that NUQLS has a better AUC-ROC value over all OoD datasets, often by a significant margin.
@@
H.3 Comparison to Sampling-LLA, VaLLA and ELLA
We now compare NUQLS against Sampling-LLA, VaLLA and ELLA. Due to issues with convergence, performance, memory usage and package compatibility when either running source code or implementing methods from instructions given in (Antorán et al., 2022) and (Ortega et al., 2023), we instead compare NUQLS against the results for Sampling-LLA, VaLLA and ELLA taken verbatim from (Ortega et al., 2023, Figure 3). We train a $2$-layer MLP, with $200$ hidden-units in each layer, on the MNIST dataset, according to the experimental details given in (Ortega et al., 2023). We then compare the accuracy, NLL, ECE, Brier score and the OOD-AUC metric of NUQLS against those reported for Sampling-LLA, VaLLA and ELLA. The results are shown in Table˜6. We display the mean and standard deviation for NUQLS; as is quoted in (Ortega et al., 2023), the standard deviation for the other methods was below $10^{-4}$ in magnitude, and was thus omitted. We see that NUQLS outperforms Sampling-LLA, VaLLA and ELLA in accuracy, NLL, ECE, and Brier score, and is within one-sixth of a standard deviation of the leading OOD-AUC value, attained by Sampling-LLA. While we cannot directly comment on differences in computation time between our method and these LLA extensions, we were able to run the source code for VaLLA for the experiments in Table˜2, where NUQLS was an order-of-magnitude faster than VaLLA in wall-time. In (Ortega et al., 2023, Figure 3 (right)), we also observe that VaLLA is an order-of-magnitude faster than both Sampling-LLA and ELLA. We also compare with (Ortega et al., 2023, Table 1), where a ResNet20 and ResNet34 model is trained on CIFAR-10. The results are displayed in Table˜7. We see that NUQLS is competitive with all other methods in this setting.
@@ H.3 Comparison to Sampling-LLA, VaLLA and ELLA
Table 7: Comparing performance of NUQLS against Sampling-LLA, VaLLA and ELLA on CIFAR-10, with ResNet20 and ResNet32. Original results taken from (Ortega et al., 2023). Purple figures correspond to the top result, while blue figures are the second-best result.
ResNet20 ResNet32
Model ACC NLL ECE ACC NLL ECE
ELLA $92.5$ $0.233$ $0.009$ $93.5$ $0.215$ $0.008$
Sampled LLA $92.5$ $0.231$ $0.006$ $93.5$ $0.217$ $0.008$
VaLLA $92.6$ $0.228$ $0.007$ $93.5$ $0.211$ $0.007$
NUQLS $92.5$ $0.228$ $0.006$ $93.4$ $0.215$ $0.007$
H.4 Comparison to BDE
Figure˜4 displays the performance of BDE on the toy regression problem from Figure˜1. We also compare BDE against NUQLS on several UCI regression tasks in Table˜8.
Figure 4: Comparison of BDE, DE and NUQLS on the toy regression problem from Figure˜1. We can see that the uncertainty of the BDE method is quite small.
Table 8: Comparing performance of NUQLS and BDE on UCI regression tasks. We see that NUQLS outperforms BDE on all tasks.
Dataset Method RMSE $\downarrow$ NLL $\downarrow$ ECE $\downarrow$
Energy BDE $416\pm0.0$390.039 $0.125\pm$$0.212$ $008\pm0.0$050.005
NUQLS $\mathbf{0.047\pm}$$\mathbf{0.006}$ $\mathbf{-2.400\pm}$$\mathbf{0.209}$ $\mathbf{0.002\pm}$$\mathbf{0.002}$
Concrete BDE $714\pm0.0$540.054 $1.563\pm$$0.449$ $063\pm0.0$120.012
NUQLS $\mathbf{0.330\pm}$$\mathbf{0.047}$ $\mathbf{-0.316\pm}$$\mathbf{0.501}$ $\mathbf{0.003\pm}$$\mathbf{0.001}$
Kin8nm BDE $851\pm0.0$370.037 $1.383\pm$$0.582$ $042\pm0.0$120.012
NUQLS $\mathbf{0.252\pm}$$\mathbf{0.005}$ $\mathbf{-0.796\pm}$$\mathbf{0.025}$ $\mathbf{0.000\pm}$$\mathbf{0.000}$
H.5 UCI Regression
We present the results for select UCI regression datasets in Table˜1 in the main body; we present results for several more datasets in Table˜9.
@@ H.5 UCI Regression
Concrete NUQLS $\mathbf{0.330\pm}$$\mathbf{0.047}$ $\mathbf{-0.316\pm}$$\mathbf{0.501}$ $\mathbf{0.003\pm}$$\mathbf{0.001}$ $\mathbf{7.339}~(\mathbf{0.185})$
DE $379\pm0.0$190.019 $\mathbf{-0.574\pm}$$\mathbf{0.098}$ $\mathbf{0.002\pm}$$\mathbf{0.002}$ $29.047$
LLA $\mathbf{0.333\pm}$$\mathbf{0.050}$ $\mathbf{-0.294\pm}$$\mathbf{0.479}$ $\mathbf{0.003\pm}$$\mathbf{0.002}$ 7.451(0.297)~7.451(0.297)
SWAG $\mathbf{0.334\pm}$$\mathbf{0.050}$ $\mathbf{-0.562\pm}$$\mathbf{0.224}$ $009\pm0.0$060.006 43.416(36.262)43.416~(36.262)
Naval NUQLS $\mathbf{0.049\pm}$$\mathbf{0.012}$ $\mathbf{-2.546\pm}$$\mathbf{0.134}$ $\mathbf{0.002\pm}$$\mathbf{0.002}$ $\mathbf{11.360}~(\mathbf{0.295})$
DE $076\pm0.0$060.006 $1.761\pm$$0.250$ $093\pm0.0$400.040 $96.570$
LLA $070\pm0.0$220.022 $25.292\pm$$17.570$ $192\pm0.0$290.029 140.724(129.659)140.724~(129.659)
SWAG $1.130\pm$$1.500$ $0.303\pm$$1.091$ $084\pm0.0$220.022 103.727(92.662)103.727~(92.662)
CCPP NUQLS $244\pm0.0$080.008 $885\pm0.0$200.020 $\mathbf{0.000\pm}$$\mathbf{0.000}$ $\mathbf{6.698}~(\mathbf{0.174})$
@@
LLA $243\pm0.0$070.007 $29.420\pm$$4.565$ $163\pm0.0$080.008 38.572(32.048)38.572~(32.048)
SWAG $252\pm0.0$120.012 $849\pm0.0$380.038 $\mathbf{0.001\pm}$$\mathbf{0.002}$ 73.357(66.833)~73.357~(66.833)
Wine NUQLS $\mathbf{0.789\pm}$$\mathbf{0.042}$ $\mathbf{0.284\pm}$$\mathbf{0.066}$ $\mathbf{0.001\pm}$$\mathbf{0.000}$ $\mathbf{1.164}~(\mathbf{0.115})$
@@
LLA $\mathbf{0.792\pm}$$\mathbf{0.041}$ $1.012\pm$$0.182$ $009\pm0.0$040.004 1.389(0.340)1.389~(0.340)
SWAG $\mathbf{0.798\pm}$$\mathbf{0.038}$ $\mathbf{0.367\pm}$$\mathbf{0.103}$ $005\pm0.0$030.003 12.856(11.807)12.856~(11.807)
Yacht NUQLS $\mathbf{0.042\pm}$$\mathbf{0.013}$ $\mathbf{-1.561\pm}$$\mathbf{2.319}$ $\mathbf{0.012\pm}$$\mathbf{0.010}$ $\mathbf{3.390}~(\mathbf{0.164})$
DE $0.647\pm$$0.121$ $2.032\pm$$0.349$ $\mathbf{0.016\pm}$$\mathbf{0.008}$ $40.132$
LLA $\mathbf{0.043\pm}$$\mathbf{0.014}$ $\mathbf{-2.733\pm}$$\mathbf{0.468}$ $\mathbf{0.011\pm}$$\mathbf{0.006}$ 3.403(0.177)3.403~(0.177)
SWAG $\mathbf{0.044\pm}$$\mathbf{0.014}$ $2.565\pm$$0.118$ $067\pm0.0$250.025 19.408(15.822)19.408~(15.822)
Song NUQLS $\mathbf{0.839\pm}$$\mathbf{0.014}$ $646\pm0.0$560.056 $1\pm0.0000.0$$0$ $\mathbf{295.058}~(\mathbf{91.673})$
DE $\mathbf{0.846\pm}$$\mathbf{0.006}$ $\mathbf{0.180\pm}$$\mathbf{0.013}$ $5\pm0.0000.0$$0$ $2562.789$
LLA-D $\mathbf{0.851\pm}$$\mathbf{0.029}$ $456\pm0.0$930.093 $\mathbf{0.000\pm}$$\mathbf{0.000}$ 413.814(210.429)413.814~(210.429)
SWAG $845\pm0.0$020.002 $680\pm0.0$620.062 $003\pm0.0$010.001 3477.825(3274.440)3477.825~(3274.440)
H.6 eNUQLS
In Figure˜5 we demonstrate the performance of an ensembled version of NUQLS, eNUQLS. This method is similar to a Mixture of Laplace Approximations (Eschenhagen et al., 2021). We observe excellent separation of the variances between correct and incorrect/OoD test groups for eNUQLS, especially for CIFAR-10 on ResNet9. Note that there is significant cost to ensembling our method, and we provide this figure simply to illustrate performance capacity.
@@
H.7 Additional VMSP Results
Figure 6: Violin plot of VMSP, for (top left) ResNet50 SVHN, (top right) ResNet50 ImageNet, (bottom left) WRN-34-1 FMNIST, and (bottom right) ResNet9 FMNIST.
We use this section to expand the empirical evaluation of NUQLS. In the top left of Figure˜6, we observe that on a ResNet50 model trained on SVHN, with CIFAR-10 as the OoD test data, NUQLS outperforms all other methods. In the top right of Figure˜6, we evaluated NUQLS on the ImageNet dataset in order to display the scalability of our method to larger datasets. We employed the pre-trained weights for a ResNet50, as found on torch.hub, and used the ImageNet-o dataset as our OOD test set (Hendrycks et al., 2021). Due to resource budget constraints, we were only able to compare NUQLS against a baseline method, though future research could include a comparison of other methods against NUQLS, using VMSP, on ImageNet. As can be seen, we see excellent separation between correct and incorrect predictions, while only adequate separation between correct and OOD points. We note from Bitterwolf et al. (2023) that approximately $21\%$ of the images in ImageNet-o are actually in-distribution; hence, we see that NUQLS correctly quantifies the variance of these test sets. We also provide comparison of NUQLS with additional competing methods. In the bottom left of Figure˜6, we compare NUQLS against the BE method. Note that BE requires modification to the structure of a neural network; hence, we compare BE on a modified WideResNet-34-1, of which an implementation of the correct modifications existed (Franchi et al., 2023), trained on FashionMNIST. We again observe that NUQLS shows greater separation than BE for VMSP across correct predictions versus incorrect or OOD predictions. Finally, in the bottom right of Figure˜6, we compare NUQLS against further competing methods, specifically BDE, SGLD (Welling & Teh, 2011) (which we include to demonstrate the inferior performance of MCMC methods in this regime) and SNGP on ResNet9 trained on FMNIST (note SNGP requires modifications to the training procedure). We see that BDE performs similarly to DE, and SLGD performs similarly to SWAG, which is unsurprising considering their respective constructions. We also observe that NUQLS still performs the strongest across all methods using the VMSP violin plots. The respective median and skew values for Figure˜6 are displayed in Table˜10 and Table˜11; these values also evidence the strong performance of NUQLS.
H.8 Image Classification Skew and Median
The median and sample skew for the VMSP in Figure˜3 and Figure˜6 is found in Table˜10 and Table˜11.
Table 10: Sample median and sample skew of variance from Figure˜3 and Figure˜6. IDC = ID correct, IDIC = ID incorrect, FMNIST = FashionMNIST. The median and skew for a method is compared against a baseline method. We expect positive skew for the ID correct (IDC) variances, and a negative skew for the ID incorrect (IDIC) and OoD variances. If the median or skew for a method is worse than the corresponding baseline, then it is written in gray.
@@ H.8 Image Classification Skew and Median
ResNet9 IDC $\downarrow$ $4.76\times 10^{-15}$ $1.31\times 10^{-5}$ $4.64\times 10^{-7}$ $5.05\times 10^{-7}$ $9.02\times 10^{-10}$ $0.020$
@@
[…]
@@
ResNet50 IDC $\downarrow$ $8.61\times 10^{-15}$ $3.59\times 10^{-6}$ $1.71\times 10^{-6}$ $2.18\times 10^{-8}$ $1.61\times 10^{-6}$ $0.020$
SVHN IDIC $\uparrow$ $0.217$ $0.0569$ $0.0302$ $\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{3.96\times 10^{-4}}$ $0.0241$ $0.020$
OoD $\uparrow$ $0.233$ $0.127$ $0.0789$ $\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{7.51\times 10^{-4}}$ $0.0252$ $0.020$
ResNet50 IDC $\downarrow$ $0.00$ - - - - $0.020$
ImageNet IDIC $\uparrow$ $0.232$ - - - - $0.020$
OoD $\uparrow$ $0.100$ - - - - $0.020$
@@
ResNet50 IDC $\uparrow$ $2.2$ $2.42$ $3.52$ $6.91$ $4.15$ $0.989$
[…]
ResNet50 IDC $\uparrow$ $1.52$ - - - - $1.21$
ImageNet IDIC $\downarrow$ 0.88-0.88 - - - - $1.10$
OoD $\downarrow$ $0.179$ - - - - $0.985$
Table 11: Sample median and sample skew of variance from Figure˜6. IDC = ID correct, IDIC = ID incorrect, FMNIST = FashionMNIST. The median and skew for a method is compared against a baseline method. We expect positive skew for the ID correct (IDC) variances, and a negative skew for the ID incorrect (IDIC) and OoD variances. If the median or skew for a method is worse than the corresponding baseline, then it is written in gray.
Median NUQLS BE BDE SGLD SNGP BASE
WRN-34-1 IDC $\downarrow$ $0.00$ $8.31\times 10^{-5}$ - - - $0.020$
FMNIST IDIC $\uparrow$ $0.178$ $0.045$ - - - $0.020$
OoD $\uparrow$ $0.173$ $0.107$ - - - $0.020$
ResNet9 IDC $\downarrow$ $4.76\times 10^{-15}$ - $3.35\times 10^{-20}$ $6.79\times 10^{-8}$ $9.19\times 10^{-7}$ $0.020$
[…]
Sample Skew
WRN-34-1 IDC $\uparrow$ $3.1$ $2.8$ - - - $1.01$
FMNIST IDIC $\downarrow$ 0.342-0.342 $0.944$ - - - $1.11$
OoD $\downarrow$ 0.0875-0.0875 0.272-0.272 - - - $1.11$
ResNet9 IDC $\uparrow$ $2.4$ - $6.46$ $5.49$ $29.8$ $0.93$
[…]
Appendix I Experiment Details